$A$ student says that if you throw a die,it will show up $1$ or not $1$. Therefore,the probability of getting $1$ and the probability of getting 'not $1$' each is equal to $\frac{1}{2}$. Is this correct? Give reasons.

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(B) No,this is not correct.
When a fair die is thrown,the total number of possible outcomes is $6$,which are ${1, 2, 3, 4, 5, 6}$.
The probability of getting $1$ is $P(1) = \frac{\text{Number of favorable outcomes}}{\text{Total number of outcomes}} = \frac{1}{6}$.
The probability of getting 'not $1$' is $P(\text{not } 1) = 1 - P(1) = 1 - \frac{1}{6} = \frac{5}{6}$.
Since $\frac{1}{6} \neq \frac{5}{6}$,the student's claim that each probability is $\frac{1}{2}$ is incorrect.

Explore More

Similar Questions

There are boards numbered $1$ to $50$ in a box. One board is selected at random. Find the probability that the number on the board is a multiple of $10$. (in $/10$)

In a packet of $500$ screws,$80$ screws are defective. Find the probability that a screw selected at random from the packet is non-defective. (in $/25$)

$A$ balanced die is rolled once. Then,the probability that the number on the die is a composite number is .......

The probability of getting two heads in the experiment of tossing an unbiased coin twice is .........

The probability of getting a bad egg in a lot of $400$ eggs is $0.035$. The number of bad eggs in the lot is:

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo